# Shadows of generalised Hayward spacetimes : in vacuum and with plasma

### v1

@@ III Null Geodesics and Shadow
$\mathit{p}_{\alpha}=\frac{1}{2}\frac{\partial\mathcal{L}}{\partial\dot{x}^{\alpha}},$ (5)
with the dot denoting derivative with respect to the affine parameter. For the static spherically symmetric metric considered, the conserved momenta $(\mathit{p}_{\phi},\mathit{p}_{t})$, corresponding to the two Killing vectors, gives $r^{2}\sin^{2}\theta\dot{\phi}\equiv L$ and $-f_{1}(r)\dot{t}\equiv-E$, respectively. The total angular momentum per unit mass and total energy per unit mass is denoted by $L$ and $E$, respectively.
@@
$0=\frac{E^{2}}{f_{1}(r)}-f(r)\left(\frac{\partial\mathcal{S}_{r}}{\partial r}\right)^{2}-\frac{1}{r^{2}}\left(\frac{\partial\mathcal{S}_{\theta}}{\partial\theta}\right)^{2}-\frac{1}{r^{2}\sin^{2}\theta}L^{2}.$ (6)
@@
$p_{\theta}^{2}$ $\equiv\left(\frac{\partial\mathcal{S}}{\partial\theta}\right)^{2}=\mathcal{K}-L^{2}\csc^{2}\theta,$ (7)
$p_{r}^{2}$ $\equiv\left(\frac{\partial\mathcal{S}}{\partial r}\right)^{2}=\frac{1}{f(r)}\left(\frac{E^{2}}{f_{1}(r)}-\frac{\mathcal{K}}{r^{2}}\right),$ (8)
with $\mathcal{K}$ being the Carter constant.Without loss of generality, we can restrict our analysis to the equatorial plane ($\theta=\frac{\pi}{2}$) due to the spherical symmetry of the spacetime. This would entail $\frac{\partial\theta}{\partial\lambda}=0$ and hence one gets $\mathcal{K}=L^{2}$ in such cases of equatorial motion, as is evident from eq.(7). The radial equation simplifies to
$\frac{1}{f(r)}\dot{r}=\sqrt{\frac{1}{f(r)f_{1}(r)}\left(E^{2}-\frac{f_{1}(r)\mathcal{K}}{r^{2}}\right)}\implies\dot{r}=L\sqrt{\frac{f(r)}{f_{1}(r)}\left(\frac{1}{b^{2}}-\frac{f_{1}(r)}{r^{2}}\right)},$ (9)
@@
$\left(\frac{p_{r}}{p_{\phi}}\right)^{2}=\left(\frac{dr}{d\phi}\right)^{2}=r^{4}\frac{f(r)}{f_{1}(r)}\left(\frac{1}{b^{2}}-V_{\rm eff}\right),$ (10)
@@
$V_{\rm eff}(r)=\frac{f_{1}(r)}{r^{2}}.$ (11)
For circular geodesics, it is known that $\dot{r}=\ddot{r}=0$. Choosing $M=1$ makes the coordinate $x$, introduced in eq.(3), equivalent to radial coordinate $r$. However, to avoid confusion we will use $x$ in the following sections when discussing equations specific to the generalised Hayward metric. Hence, from eq.(8) and its derivative, we get the following set of equations,
$\dot{x}$ $=\frac{f(x)}{f_{1}(x)}\left(\frac{1}{b^{2}}-\frac{f_{1}(x)}{x^{2}}\right)=0,$ (12)
$\ddot{x}$ $=\frac{\left(f_{1}(x)f^{\prime}(x)-f(r)f_{1}^{\prime}(x)\right)}{f_{1}^{2}(x)}\left(\frac{1}{b^{2}}-\frac{f_{1}(x)}{x^{2}}\right)+\frac{f(x)}{x^{2}f_{1}(x)}\left(-f_{1}^{\prime}(x)+\frac{2f_{1}(x)}{x}\right)=0,$ (13)
solving which enables us to determine the photon sphere radius $x_{\rm{ph}}$ and the corresponding impact parameter $b$. The critical orbits are unstable if $\dddot{r}>0$ and stable if $\dddot{r}<0$. Note that the unstable null circular orbits are the photon spheres and the stable null circular orbits are known as anti-photon spheres. The shadow caste by these circular null geodesics as observed by an observer located at $x_{o}$ is given as
$x_{sh}^{2}=\frac{b^{2}}{\frac{f(x_{o})}{f_{1}(x_{o})}\left(1-\frac{f_{1}(x_{o})}{x_{o}^{2}}b^{2}\right)}.$ (14)
@@
III.1 Shadows with plasma medium
@@
$n^{2}(r)=1-\frac{\omega_{p}^{2}(r)}{\omega^{2}}.$ (16)
In presence of surrounding plasma, the null geodesic equation of motion, given in eq.(6), gets modified to
@@
$p_{\theta}^{2}$ $\equiv\left(\frac{\partial\mathcal{S}}{\partial\theta}\right)^{2}=\mathcal{K}-L^{2}\csc^{2}\theta$ (17)
$p_{r}^{2}$ $\equiv\left(\frac{\partial\mathcal{S}}{\partial r}\right)^{2}=\frac{1}{f(r)}\left(\frac{E^{2}}{f_{1}(r)}-\frac{\mathcal{K}}{r^{2}}-\omega_{p}^{2}(r)\right).$ (18)
@@
$\omega(r)=\frac{\omega_{0}}{\sqrt{f_{1}(r)}},$ (19)
@@
$n^{2}=1-\frac{\omega_{p}^{2}}{\frac{\omega_{0}^{2}}{f_{1}(r)}}.$ (20)
@@
Similar to the vacuum case, we translate to $x$ coordinate and also introduce an new parameter $\Omega(x)=\frac{\omega_{p}^{2}}{E^{2}}$. Hence, the equivalent condition to eq.(12) and (13) for the circular orbits in presence of plasma is,
$\dot{x}$ $=\frac{f(x)}{f_{1}(x)}\left(\frac{1}{b^{2}}-\frac{f_{1}(x)}{x^{2}}-\frac{\Omega(x)f_{1}(x)}{b^{2}}\right)=0$ (21)
$\ddot{x}$ $=\left(\frac{f(x)}{f_{1}(x)}\right)^{\prime}\left(\frac{1}{b^{2}}-\frac{f_{1}(x)}{x^{2}}-\frac{\Omega(x)f_{1}(x)}{b^{2}}\right)+\frac{f(x)}{f_{1}(x)}\left(-\frac{f_{1}^{\prime}(x)}{x^{2}}+\frac{2f_{1}(x)}{x^{3}}-\frac{(\Omega(x)f_{1}(x)^{\prime})}{b^{2}}\right)=0$ (22)
@@
$b^{2}=\frac{x^{2}}{f_{1}}\left(1-\Omega f_{1}\right)\Big{|}_{x=x_{\rm ph}}.$ (23)
@@
$f(x)=0,$ (24)
$\frac{1}{f_{1}(x)}(1-\Omega(x)f_{1}(x))\Big{(}-f_{1}^{\prime}(x)+\frac{2f_{1}(x)}{x}\Big{)}-(\Omega(x)f_{1}(x))^{\prime}=0.$ (25)
@@
$x_{\rm sh}^{2}$ $=\frac{\frac{x_{\rm ph}^{2}}{f_{1}(x_{\rm ph})}\left(1-\Omega(x_{\rm ph})f_{1}(x_{\rm ph})\right)}{\frac{f(x_{o})}{f_{1}(x_{o})}\left(1-\frac{b^{2}}{x^{2}_{o}}f_{1}(x_{o})-\Omega(x_{o})f_{1}(x_{o})\right)}.$ (26)
@@
$x_{\rm sh}^{2}\Big{|}_{x_{o}\rightarrow\infty}=\frac{\frac{x_{\rm ph}^{2}}{f_{1}(x_{\rm ph})}\left(1-\Omega(x_{\rm ph})f_{1}(x_{\rm ph})\right)}{\left(1-\Omega(\infty)\right)}.$ (27)
Note that the numerator in the above equation is required to be positive as it represents the impact parameter, as shown in eq.(23), which is a physical quantity and hence cannot be negative.
Now let us apply the above results to the two plasma profiles we are interested. We begin by considering the homogeneous plasma profile given by $\Omega=k_{0}$. Solving equations (24) and (25) with $\Omega=k_{0}$ gives the impact parameter
$b^{2}=\frac{x_{\rm ph}^{2}}{f_{1}(x_{\rm ph})}\left(1-k_{0}f_{1}(x_{\rm ph})\right)$ (28)
with the shadow radius obtained from eq.(27)
$x_{\rm sh}^{2}=\frac{\frac{x_{\rm ph}^{2}}{f_{1}(x_{\rm ph})}\left(1-k_{0}f_{1}(x_{\rm ph})\right)}{\left(1-k_{0}\right)}.$ (29)
Since the impact parameter, given by eq.(28), is a physical quantity which needs to be positive, the denominator of eq.(29) also needs to be positive, hence giving $10\leq\kappa_{0}<$ 1. Further constraints on this parameter would depend on the refractive index condition and will be studied for each spacetime from the generalised Hayward metric.
For the non-homogeneous case, $\Omega=\frac{k_{x}}{x}$. Similar to the homogeneous profile, solving equations (24) and (25) we get the impact parameter to be
$b^{2}=\frac{x_{\rm ph}^{2}}{f_{1}(x_{\rm ph})}\left(1-\frac{k_{x}}{x_{\rm ph}}f_{1}(x_{\rm ph})\right).$ (30)
Note that as $x\rightarrow\infty$, the plasma profile $\Omega\rightarrow 0$, hence the shadow radius in eq.(27) reduces to
$x_{\rm sh}^{2}=b^{2}=\frac{x_{\rm ph}^{2}}{f_{1}(x_{\rm ph})}\left(1-\frac{k_{x}}{x_{\rm ph}}f_{1}(x_{\rm ph})\right).$ (31)
@@
IV Shadow of Generalised Hayward metric
In this section, we will apply the above results for the various spacetimes of the generalized Hayward metric given in eq.(3). and summarized in Table 1.
IV.1 Schwarzschild black hole $\sigma=1$, $\kappa=0$:
@@
By setting $\sigma=1$, $\kappa=0$ in eq.(3), we recover the standard Schwarzschild solution. The shadow radius of Schwarzschild BH was first analyzed by [36]. This is one of the most well-known results in the literature, and we include these results for completeness. Solving eqns.(12) and (13), we get $x_{\rm ph}=3$, $b^{2}=27$ and the shadow radius being $x_{\rm sh}=3\sqrt{3}$.
We now focus on the influence of the plasma on the shadow of Schwarzschild BH. For the case of homogeneous plasma, the eq.(24) does not lead to any solutions. However, eq.(25) reduces to
$(1-k_{0})x^{2}+(4k_{0}-3)x-4k_{0}=0.$ (32)
Solving this quadratic equation we find that the $x_{\rm ph}>x_{\rm h}$ for all the range of $k_{0}$ where $x_{\rm h}$ is the horizon radius. Note that the shadow radius diverges at $k_{0}\rightarrow 1$, which is observed in the Fig.2(a).
Similarly, for the case of non-homogeneous plasma, the eq.(25) reduces to
$02x^{3}+(-k_{x}-6)x^{2}+4k_{x}x-4k_{x}=$ 0 . (33)
Substituting the solution of the above equation in eq.(31), it can be checked that the shadow radius becomes zero at $k_{x}=8$, beyond which the photon sphere does not exist. The shadow radius decreases monotonically as a function of the plasma parameter $k_{x}$ as shown in Fig.2(b).
IV.2 Schwarzschild Wormhole $\sigma=0,\kappa=0$:
@@
For $\sigma=\kappa=0$, the metric functions in eq.(3) become $f_{1}(x)=1$ and $f(x)=1-\frac{2}{x}$. Solving eq.(12),(13) we get $x_{\rm ph}=2,\,b^{2}=x_{\rm ph}^{2}=2,\,x_{\rm sh}=\sqrt{b^{2}}=2$.
Now we look at the influence of plasma on the shadow of Schwarzschild wormhole spacetime. For the case of homogeneous plasma profile, we observe that the photon sphere is still located at the throat $x=2$ obtained from solving $f(x)=0$. Since $f_{1}(x)=1$, the shadow radius from eq.(29) becomes independent of the plasma parameter $k_{0}$. Hence it is identical to the vacuum case.
Moving on to the case of non-homogeneous plasma profile, the photon sphere equations (24) and (25) provide us with the solutions $x_{\rm ph}=2$ and $x_{\rm ph}=2-k_{x}$. The shadow radius corresponding to the two photon radius is thus $x_{\rm sh}^{2}=2(2-k_{x})$ and $x_{\rm sh}^{2}=-\frac{k_{x}}{4}$ respectively. Of the above solutions, only the first one is physically reasonable. Hence, from $x_{\rm sh}^{2}=2(2-k_{x})$, we can deduce that the shadow radius decreases monotonically with increasing plasma parameter $k_{x}$. At $k_{x}=2$, the shadow radius vanishes and becomes unphysical beyond this point as shown in Fig.3.
@@
IV.3 Damour-Solodukhin Wormhole $00<\sigma<1,\kappa=$ 0 :
For the Damour-Solodukhin wormhole, the metric components simplify to $f_{1}(x)=1-\frac{2\sigma}{x}$ and $f(x)=1-\frac{2}{x}$. Unlike the previous cases, for this wormhole we have two possible solutions to eq.(12), which will be discussed in detail one by one. The first set of solutions corresponds to $f(x)=0$, which gives $x_{\rm ph}=2$. The critical impact parameter for this is obtained by solving eq.(13), which gives $b^{2}=\frac{4}{1-\sigma}$. It is evident that these circular geodesics are located at the throat. Note that the photons will be observable to an asymptotic observer only if the photon orbits are unstable which is determined by the condition $\dddot{x}|_{x=x_{\rm ph}}>0$. Hence we check the stability of the null geodesics by computing $\dddot{x}|_{x=x_{\rm ph}}$ which for this first set of solutions is given by
$\dddot{x}|_{x=x_{\rm ph}}=\frac{2-3\sigma}{8(1-\sigma)}.$ (34)

### v2

@@ 3 Null Geodesics and Shadow
$\mathit{p}_{\alpha}=\frac{1}{2}\frac{\partial\mathcal{L}}{\partial\dot{x}^{\alpha}},$ (3.4)
with the dot denoting derivative with respect to the affine parameter. For the static spherically symmetric metric considered, the conserved momenta $(\mathit{p}_{\phi},\mathit{p}_{t})$, corresponding to the two Killing vectors, gives $r^{2}\sin^{2}\theta\dot{\phi}\equiv L$ and $f_{1}(r)\dot{t}\equiv-E$, respectively. The total angular momentum per unit mass and total energy per unit mass is denoted by $L$ and $E$, respectively.
@@
$0=\frac{E^{2}}{f_{1}(r)}-f(r)\left(\frac{\partial\mathcal{S}_{r}}{\partial r}\right)^{2}-\frac{1}{r^{2}}\left(\frac{\partial\mathcal{S}_{\theta}}{\partial\theta}\right)^{2}-\frac{1}{r^{2}\sin^{2}\theta}L^{2}.$ (3.5)
@@
$p_{\theta}^{2}$ $\equiv\left(\frac{\partial\mathcal{S}}{\partial\theta}\right)^{2}=\mathcal{K}-L^{2}\csc^{2}\theta,$ (3.6)
$p_{r}^{2}$ $\equiv\left(\frac{\partial\mathcal{S}}{\partial r}\right)^{2}=\frac{1}{f(r)}\left(\frac{E^{2}}{f_{1}(r)}-\frac{\mathcal{K}}{r^{2}}\right),$ (3.7)
with $\mathcal{K}$ being the Carter constant.Without loss of generality, we can restrict our analysis to the equatorial plane ($\theta=\frac{\pi}{2}$) due to the spherical symmetry of the spacetime. This would entail $\frac{\partial\theta}{\partial\lambda}=0$ and hence one gets $\mathcal{K}=L^{2}$ in such cases of equatorial motion, as is evident from eq.(3.6). The radial equation simplifies to
$\frac{1}{f(r)}\dot{r}=\sqrt{\frac{1}{f(r)f_{1}(r)}\left(E^{2}-\frac{f_{1}(r)\mathcal{K}}{r^{2}}\right)}\implies\dot{r}=L\sqrt{\frac{f(r)}{f_{1}(r)}\left(\frac{1}{b^{2}}-\frac{f_{1}(r)}{r^{2}}\right)},$ (3.8)
@@
$\left(\frac{p_{r}}{p_{\phi}}\right)^{2}=\left(\frac{dr}{d\phi}\right)^{2}=r^{4}\frac{f(r)}{f_{1}(r)}\left(\frac{1}{b^{2}}-V_{\rm eff}\right),$ (3.9)
@@
$V_{\rm eff}(r)=\frac{f_{1}(r)}{r^{2}}.$ (3.10)
For circular geodesics, it is known that $\dot{r}=\ddot{r}=0$. Choosing $M=1$ makes the coordinate $x$, introduced in eq.(2.3), equivalent to radial coordinate $r$. However, to avoid confusion we will use $x$ in the following sections when discussing equations specific to the generalised Hayward metric. Hence, from eq.(3.7) and its derivative, we get the following set of equations,
$\dot{x}$ $=\frac{f(x)}{f_{1}(x)}\left(\frac{1}{b^{2}}-\frac{f_{1}(x)}{x^{2}}\right)=0,$ (3.11)
$\ddot{x}$ $=\frac{\left(f_{1}(x)f^{\prime}(x)-f(r)f_{1}^{\prime}(x)\right)}{f_{1}^{2}(x)}\left(\frac{1}{b^{2}}-\frac{f_{1}(x)}{x^{2}}\right)+\frac{f(x)}{x^{2}f_{1}(x)}\left(-f_{1}^{\prime}(x)+\frac{2f_{1}(x)}{x}\right)=0,$ (3.12)
solving which enables us to determine the photon sphere radius $x_{\rm{ph}}$ and the corresponding impact parameter $b$. The critical orbits are unstable if $\dddot{x}>0$ and stable if $\dddot{x}<0$. Note that the unstable null circular orbits are the photon spheres and the stable null circular orbits are known as anti-photon spheres.
Using the above information and plugging them into eq.(3.1) and eq.(3.2) we obtain:
$x_{sh}^{2}=\frac{b^{2}}{\frac{f(x_{o})}{f_{1}(x_{o})}\left(1-\frac{f_{1}(x_{o})}{x_{o}^{2}}b^{2}\right)}.$ (3.13)
Note that we have defined $x_{sh}=\frac{r_{sh}}{M}$. However as emphasised earlier we are setting $M=1$, which makes both these definitions equivalent.
@@
3.1 Shadows with plasma medium
@@
$n^{2}(r)=1-\frac{\omega_{p}^{2}(r)}{\omega^{2}}.$ (3.15)
In the presence of surrounding plasma, the null geodesic equation of motion, given in eq.(3.5), gets modified to
@@
$p_{\theta}^{2}$ $\equiv\left(\frac{\partial\mathcal{S}}{\partial\theta}\right)^{2}=\mathcal{K}-L^{2}\csc^{2}\theta$ (3.16)
$p_{r}^{2}$ $\equiv\left(\frac{\partial\mathcal{S}}{\partial r}\right)^{2}=\frac{1}{f(r)}\left(\frac{E^{2}}{f_{1}(r)}-\frac{\mathcal{K}}{r^{2}}-\omega_{p}^{2}(r)\right).$ (3.17)
@@
$\omega(r)=\frac{\omega_{0}}{\sqrt{f_{1}(r)}},$ (3.18)
@@
$n^{2}=1-\frac{\omega_{p}^{2}}{\frac{\omega_{0}^{2}}{f_{1}(r)}}.$ (3.19)
@@
Similar to the vacuum case, we translate to $x$ coordinate and also introduce an new parameter $\Omega(x)=\frac{\omega_{p}^{2}}{E^{2}}$. Hence, the equivalent condition to eq.(3.11) and (3.12) for the circular orbits in presence of plasma is,
$\dot{x}$ $=\frac{f(x)}{f_{1}(x)}\left(\frac{1}{b^{2}}-\frac{f_{1}(x)}{x^{2}}-\frac{\Omega(x)f_{1}(x)}{b^{2}}\right)=0$ (3.20)
$\ddot{x}$ $=\left(\frac{f(x)}{f_{1}(x)}\right)^{\prime}\left(\frac{1}{b^{2}}-\frac{f_{1}(x)}{x^{2}}-\frac{\Omega(x)f_{1}(x)}{b^{2}}\right)+\frac{f(x)}{f_{1}(x)}\left(-\frac{f_{1}^{\prime}(x)}{x^{2}}+\frac{2f_{1}(x)}{x^{3}}-\frac{(\Omega(x)f_{1}(x)^{\prime})}{b^{2}}\right)=0$ (3.21)
@@
$b^{2}=\frac{x^{2}}{f_{1}}\left(1-\Omega f_{1}\right)\Big{|}_{x=x_{\rm ph}}.$ (3.22)
@@
$f(x)=0,$ (3.23)
$\frac{1}{f_{1}(x)}(1-\Omega(x)f_{1}(x))\Big{(}-f_{1}^{\prime}(x)+\frac{2f_{1}(x)}{x}\Big{)}-(\Omega(x)f_{1}(x))^{\prime}=0.$ (3.24)
@@
$x_{\rm sh}^{2}$ $=\frac{\frac{x_{\rm ph}^{2}}{f_{1}(x_{\rm ph})}\left(1-\Omega(x_{\rm ph})f_{1}(x_{\rm ph})\right)}{\frac{f(x_{o})}{f_{1}(x_{o})}\left(1-\frac{b^{2}}{x^{2}_{o}}f_{1}(x_{o})-\Omega(x_{o})f_{1}(x_{o})\right)}.$ (3.25)
@@
$x_{\rm sh}^{2}\Big{|}_{x_{o}\rightarrow\infty}=\frac{\frac{x_{\rm ph}^{2}}{f_{1}(x_{\rm ph})}\left(1-\Omega(x_{\rm ph})f_{1}(x_{\rm ph})\right)}{\left(1-\Omega(\infty)\right)}.$ (3.26)
Note that the numerator in the above equation is required to be positive as it represents the impact parameter, as shown in eq.(3.22), which is a physical quantity and hence cannot be negative.
Now let us apply the above results to the two plasma profiles we are interested. We begin by considering the homogeneous plasma profile given by $\Omega=k_{0}$. Solving equations (3.23) and (3.24) with $\Omega=k_{0}$ gives the impact parameter
$b^{2}=\frac{x_{\rm ph}^{2}}{f_{1}(x_{\rm ph})}\left(1-k_{0}f_{1}(x_{\rm ph})\right)$ (3.27)
with the shadow radius obtained from eq.(3.26)
$x_{\rm sh}^{2}=\frac{\frac{x_{\rm ph}^{2}}{f_{1}(x_{\rm ph})}\left(1-k_{0}f_{1}(x_{\rm ph})\right)}{\left(1-k_{0}\right)}.$ (3.28)
Since the impact parameter, given by eq.(3.27), is a physical quantity which needs to be positive, the denominator of eq.(3.28) also needs to be positive, hence giving $0\leq\kappa_{0}<1$. Further constraints on this parameter would depend on the refractive index condition and will be studied for each spacetime from the generalised Hayward metric.
For the non-homogeneous case, $\Omega=\frac{k_{x}}{x}$. Similar to the homogeneous profile, solving equations (3.23) and (3.24) we get the impact parameter to be
$b^{2}=\frac{x_{\rm ph}^{2}}{f_{1}(x_{\rm ph})}\left(1-\frac{k_{x}}{x_{\rm ph}}f_{1}(x_{\rm ph})\right).$ (3.29)
Note that as $x\rightarrow\infty$, the plasma profile $\Omega\rightarrow 0$, hence the shadow radius in eq.(3.26) reduces to
$x_{\rm sh}^{2}=b^{2}=\frac{x_{\rm ph}^{2}}{f_{1}(x_{\rm ph})}\left(1-\frac{k_{x}}{x_{\rm ph}}f_{1}(x_{\rm ph})\right).$ (3.30)
@@
4 Shadow of Generalised Hayward metric
In this section, we will apply the above results for the various spacetimes of the generalized Hayward metric given in eq.(2.3). and summarized in Table 1.
4.1 Schwarzschild black hole $\sigma=1$, $\kappa=0$:
@@
By setting $\sigma=1$, $\kappa=0$ in eq.(2.3), we recover the standard Schwarzschild solution. The shadow radius of Schwarzschild BH was first analyzed by [35]. This is one of the most well-known results in the literature, and we include these results for completeness. Solving eqns.(3.11) and (3.12), we get $x_{\rm ph}=3$, $b^{2}=27$ and the shadow radius being $x_{\rm sh}=3\sqrt{3}$.
We now focus on the influence of the plasma on the shadow of Schwarzschild BH. For the case of homogeneous plasma, the eq.(3.23) does not lead to any solutions. However, eq.(3.24) reduces to
$(1-k_{0})x^{2}+(4k_{0}-3)x-4k_{0}=0.$ (4.1)
Solving this quadratic equation we find that the $x_{\rm ph}>x_{\rm h}$ for all the range of $k_{0}$ where $x_{\rm h}$ is the horizon radius. Note that the shadow radius diverges at $k_{0}\rightarrow 1$, which is observed in the Fig.(2(a)).
Similarly, for the case of non-homogeneous plasma, the eq.(3.24) reduces to
$2x^{3}+(-k_{x}-6)x^{2}+4k_{x}x-4k_{x}=0.$ (4.2)
Substituting the solution of the above equation in eq.(3.30), it can be checked that the shadow radius becomes zero at $k_{x}=8$, beyond which the photon sphere does not exist. The shadow radius decreases monotonically as a function of the plasma parameter $k_{x}$ as shown in Fig.(2(b)).
4.2 Schwarzschild Wormhole $\sigma=0,\kappa=0$:
@@
For $\sigma=\kappa=0$, the metric functions in eq.(2.3) become $f_{1}(x)=1$ and $f(x)=1-\frac{2}{x}$. Solving eq.(3.11),(3.12) we get $x_{\rm ph}=2,\,b^{2}=x_{\rm ph}^{2}=2,\,x_{\rm sh}=\sqrt{b^{2}}=2$.
Now we look at the influence of plasma on the shadow of Schwarzschild wormhole spacetime. For the case of homogeneous plasma profile, we observe that the photon sphere is still located at the throat $x=2$ obtained from solving $f(x)=0$. Since $f_{1}(x)=1$, the shadow radius from eq.(3.28) becomes independent of the plasma parameter $k_{0}$. Hence it is identical to the vacuum case.
Moving on to the case of non-homogeneous plasma profile, the photon sphere equations (3.23) and (3.24) provide us with the solutions $x_{\rm ph}=2$ and $x_{\rm ph}=2-k_{x}$. The shadow radius corresponding to the two photon radius is thus $x_{\rm sh}^{2}=2(2-k_{x})$ and $x_{\rm sh}^{2}=-\frac{k_{x}}{4}$ respectively. Of the above solutions, only the first one is physically reasonable. Hence, from $x_{\rm sh}^{2}=2(2-k_{x})$, we can deduce that the shadow radius decreases monotonically with increasing plasma parameter $k_{x}$. At $k_{x}=2$, the shadow radius vanishes and becomes unphysical beyond this point as shown in Fig.(3).
@@
4.3 Damour-Solodukhin Wormhole $0<\sigma<1,\kappa=0$ :
For the Damour-Solodukhin wormhole, the metric components simplify to $f_{1}(x)=1-\frac{2\sigma}{x}$ and $f(x)=1-\frac{2}{x}$. Unlike the previous cases, for this wormhole we have two possible solutions to eq.(3.11), which will be discussed in detail one by one. The first set of solutions corresponds to $f(x)=0$, which gives $x_{\rm ph}=2$. The critical impact parameter for this is obtained by solving eq.(3.12), which gives $b^{2}=\frac{4}{1-\sigma}$. It is evident that these circular geodesics are located at the throat. Note that the photons will be observable to an asymptotic observer only if the photon orbits are unstable which is determined by the condition $\dddot{x}|_{x=x_{\rm ph}}>0$. Hence we check the stability of the null geodesics by computing $\dddot{x}|_{x=x_{\rm ph}}$ which for this first set of solutions is given by
$\dddot{x}|_{x=x_{\rm ph}}=\frac{2-3\sigma}{8(1-\sigma)}.$ (4.3)